y=-2x-1
y=3x-16
-2x-1=3x-16
-1=5x-16
15=5x
3=x
y=-2(3)-1
y=-6-5
y=-11

Answer by stanbon(48568)

y=-2x-1
y=3x-16
-2x-1=3x-16
-1=5x-16
15=5x
3=x
y=-2(3)-1
y=-6-1
y=-7
Cheers,
Stan H.